# Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D. Salim1

### v1

@@ 3.3 Task-Aware Tuning
As shown in Figure 2 (b), the Task-Aware Tuning stage employs a classification model to align the multimodal sensor-text embeddings with human activity labels. This stage further integrates multi-channel sensor readings with activity labels for HAR. The input sensor data $\mathbf{X}$ is divided into segments of window size $L$ with 50% overlap (Li et al., 2018), forming segments $\mathbf{X}_{S}\in\mathbb{R}^{S\times C\times L}$, where $S$ is the number of segments and $C$ is the number of channels. The pretrained alignment module from first stage is used to map sensor data to activity labels, learning the relationship between temporal patterns and human activities while preserving inter-channel dependencies.
@@
Input Embedding. For each sensor channel $c$, we first retrieve the combined sensor embeddings $\hat{o}_{s}^{c}$. These embeddings are then concatenated across all sensor channels, along with their corresponding statistical information (mean and variance) as prompts, to form the final input embedding $\hat{z}=\hat{o}_{s}^{1}\oplus\hat{o}_{s}^{2}\oplus\dots\oplus\hat{o}_{s}^{C}\oplus\hat{z}_{\text{stat}}$, where $\hat{z}_{\text{stat}}$ represents the statistical information of each channel, and $C$ is the total number of channels. This ensures the model integrates both temporal and statistical characteristics of the sensor data, providing a comprehensive input for HAR.
@@
Loss Function. The input token sequence is passed through the LLM to extract a latent representation $\mathbf{H}\in\mathbb{R}^{K\times D}$, where $K$ is the number of tokens and $D$ is the embedding dimension. Due to causal masking, the model pools the final hidden state by selecting the representation of the last token, $\mathbf{h}=\mathbf{H}_{K}$, which encodes information from all preceding tokens. This pooled vector, $\mathbf{h}$, is then passed through a fully connected layer, which maps it to a vector of size $M$, where $M$ is the number of activity classes. The softmax function is applied to obtain the predicted class probabilities $\hat{y}_{i}$, which represent the likelihood of each activity class. The model then is optimized using the cross-entropy loss, defined as: $\mathcal{L}_{cls}=-\sum_{i=0}^{M-1}y_{i}\log\hat{y}_{i}$, where $y_{i}$ is the ground truth label.
4 Datasets
Figure 3: Proportion of human activity labels in each training dataset.
To evaluate the effectiveness of SensorLLM, we utilized four publicly available HAR datasets.
USC Human Activity Dataset (USC-HAD). USC-HAD (Zhang & Sawchuk, 2012) consists of six sensor readings from body-worn 3-axis accelerometers and gyroscopes, collected from 14 subjects. The data is sampled at 100 Hz across six channels and includes 12 activity class labels. For evaluation, we use data from subjects 13 and 14 as the test set, while the remaining subjects’ data are used for training. A window size $w\in[5,200]$ is used in alignment stage, and $w=200$ is used in HAR.
UCI Human Activity Recognition Dataset (UCI-HAR). UCI-HAR (Anguita et al., 2013) includes data collected from 30 volunteers performing six activities while wearing a smartphone on their waist. The embedded accelerometer and gyroscope sensors sampled data at 50 Hz across six channels. The dataset was partitioned into 70% for training and 30% for testing. A window size $w\in[5,200]$ is used in alignment stage, and $w=128$ is used in HAR.
Physical Activity Monitoring Dataset (PAMAP2). PAMAP2 (Reiss & Stricker, 2012) includes data from nine subjects wearing IMUs on their chest, hands, and ankles. IMUs capture the acceleration, gyroscope, and magnetometer data across 27 channels and include 12 activity class labels. For our experiments, data from subjects 105 and 106 are used as the test set, with the remaining subjects’ data used for training. The sample rate is downsampled from 100 Hz to 50 Hz. A window size $w\in[5,100]$ is used in alignment stage, and $w=100$ in HAR.
Mobile Health Dataset (MHealth). MHealth (Baños et al., 2014) contains body motion and vital sign recordings from ten volunteers. Sensors were placed on the chest, right wrist, and left ankle of each subject. For our experiments, we used acceleration data from the chest, left ankle, and right lower arm, along with gyroscope data from the left ankle and right lower arm, resulting in a total of 15 channels. The data is sampled at 50 Hz and includes 12 activity class labels. Data from subjects 1, 3, and 6 is used as the test set, while the remaining subjects’ data are used for training. We use a window size $w\in[5,100]$ in alignment stage and $w=100$ in HAR.
The activities associated with each dataset are shown in Figure 3.
@@ 5 Experiments
In this section we investigate the effectiveness of SensorLLM in enabling LLM to interpret and reason across temporal-numerical (i.e., time-series) data, and in performing HAR tasks. All experiments were conducted on NVIDIA A100-80G GPUs. To ensure a fair evaluation, we used the same training and testing subjects across both stages. This setup guarantees that the test data used in Task-Aware Tuning Stage was not exposed during the alignment training in Sensor-Language Alignment Stage, ensuring that the task-aware results reflect the LLM’s true ability to learn and interpret the semantic information from the input data. We chose Chronos as the TS embedder for our experiments, as it has not been pre-trained on any sensor data, making it an ideal candidate for evaluating the robustness of our approach in learning from raw sensor inputs.
@@
Metric USC-HAD UCI-HAR PAMAP2 MHealth
GPT-4o Ours GPT-4o Ours GPT-4o Ours GPT-4o Ours
BLEU-1 41.43 57.68 37.97 56.78 46.35 60.20 49.97 61.38
ROUGE-1 54.92 68.32 51.24 67.63 58.08 69.92 61.11 71.20
ROUGE-L 49.00 64.17 44.88 63.05 50.30 66.25 51.99 67.83
METEOR 30.51 45.95 26.93 45.81 37.17 52.21 38.50 51.73
SBERT 77.22 86.09 76.05 85.01 82.71 87.31 83.15 86.66
SimCSE 86.96 93.09 90.23 92.51 89.64 93.82 92.10 93.38
GPT-4o 1.67 3.11 1.61 3.2 1.90 3.77 1.69 3.69
Human 2.1 4.16 1.94 4.04 2.38 4.7 1.74 4.56
@@ 5.1 Sensor Data Understanding
Setups. For all four datasets, we trained with the following parameters in the Sensor-Language Alignment Stage: learning rate of 2e-3, 8 epochs, batch size of 4, gradient accumulation steps of 8, and maximum sequence length of 4096.
@@
Metrics. We evaluate the effectiveness of SensorLLM in understanding and generating trend descriptions from sensor data, compared to the robust GPT-4o111gpt-4o-2024-08-06(OpenAI, 2024). This evaluation primarily focuses on assessing the success of our Sensor-Language Alignment. Using a predefined prompt A.3, GPT-4o generates answers based on the input sensor data, following our text template. To comprehensively assess performance, we employed three different evaluation methods.
@@
  • NLP Metric Evaluation: We use standard NLP metrics including BLEU-1 (Papineni et al., 2002), ROUGE-1, ROUGE-L (Lin, 2004), and METEOR (Banerjee & Lavie, 2005) to measure surface-level similarity and n-gram overlap. Besides, we also report Sentent-BERT (SBERT) (Reimers & Gurevych, 2019) and SimCSE (Gao et al., 2021), which compute the semantic similarity between sentence embeddings of the model output and ground truth. For a fair comparison, we used SimCSE’s tokenizer across all metrics, except for SBERT, to better handle numerical content like timestamps.
  • GPT-4o Evaluation: We task GPT-4o with evaluating the generated text based on its alignment with ground truth. GPT-4o assessed the correctness of the trend descriptions and rated each model’s output on a scale of 1 to 5 (with 5 being the highest) with accompanying explanations based on predefined criteria. As one of the most advanced LLMs, GPT-4o’s semantic evaluation ensures a more accurate assessment of how well the models capture trend-related information.
  • Human Evaluation: Five human experts were tasked with evaluating the generated trend descriptions. Using the same scoring criteria as GPT-4o, they assessed the correctness of the descriptions, providing a human-centred evaluation of the model outputs.
A detailed introduction to the metrics and scoring criteria is provided in Appendix A.5. To ensure comprehensive evaluation, we randomly selected 200 data samples from each dataset for both SensorLLM and GPT-4o to generate descriptive and summary text. The evaluation results were averaged and reported. Due to the length of many descriptive texts, we selected 20 samples from each dataset with sequence lengths of 50 time steps or fewer for human evaluation.
@@
Results. Table 1 compares SensorLLM and GPT-4o on the Sensor Data Trend Analysis tasks, showing that our model consistently outperforms GPT-4o across all metrics. BLEU-1, ROUGE-1, ROUGE-L, and METEOR primarily focus on surface-level lexical or n-gram overlaps and SBERT and SimCSE can capture factual correctness or deeper semantic similarities. Results from all metrics confirm SensorLLM’s superior ability to generate trend descriptions that are closely aligned with the ground truth. The GPT-4o evaluation offers a more nuanced assessment, examining the semantic depth, detail, and coherence of the generated descriptions. These results illustrate that SensorLLM has a stronger capacity to interpret and process sensor data, whereas even GPT-4o struggle with tasks involving numerical data and complex trend observations (Yehudai et al., 2024). Human evaluations also show higher accuracy for SensorLLM when processing shorter input sequences. Overall, these results demonstrate the effectiveness of our Sensor-Language Alignment method, validating its ability to enhance LLMs’ understanding of sensor data. Appendix A.8 shows the examples generated by GPT-4o and SensorLLM.
(a) USC-HAD
(b) UCI-HAR
(c) PAMAP2
(d) MHealth
Figure 4: Experimental results of Task-Aware Tuning Stage from five random runs. The red dashed line indicates the highest-performing result across the experiments.
@@ 5.2 Human Activity Recognition
Setups. In this section, we evaluate the performance of SensorLLM on HAR tasks. For all datasets, experiments were run five times with the following parameters: 8 training epochs, a batch size of 4, gradient accumulation steps of 8, and a maximum sequence length of 4096. We report the macro F1 score A.7 to account for class imbalance across different activity categories.
Baselines. To benchmark our model, we compare it against two sets of baselines. The first set includes state-of-the-art time-series (TS) models: Transformer (Vaswani et al., 2017), Informer (Zhou et al., 2021), NS-Transformer (Liu et al., 2022), PatchTST (Nie et al., 2023), TimesNet (Wu et al., 2023a), and iTransformer (Liu et al., 2024c). The second set consists of state-of-the-art HAR models: DeepConvLSTM (Ordóñez & Roggen, 2016), DeepConvLSTM + Attention (DeepConvLSTMAttn) (Murahari & Plötz, 2018), and Attend (Abedin et al., 2021). Detailed introduction of the baselines can be found in the Appendix A.6.
Results. As shown in Figure 4, SensorLLM outperforms the TS baselines across all datasets and surpasses the HAR baselines on USC-HAD, MHealth, and PAMAP2.
On USC-HAD, SensorLLM improves the median performance of the top TS baseline NS-Transformer by 8% and matches the top HAR baseline Attend, with a 3% gain in the best results. For UCI-HAR, SensorLLM beats NS-Transformer by 3% in median performance but trails 2% short of Attend. On PAMAP2, SensorLLM surpasses PatchTST by 4% in median performance and exceeds Attend by 1%, with a 2% improvement in best results. For MHealth, SensorLLM outperforms both PatchTST and iTransformer by 9% and improves on Attend by 6%, with a 10% gain in best results.
The fact that the TS embedder Chronos was not pre-trained on sensor data further highlights SensorLLM’s potential to understand sensor data and achieve superior performance in HAR and other related tasks.
@@ 6 Ablation Studies
In this section, we present our ablation studies to further demonstrate the effectiveness of our proposed method.
Figure 5: Experiments demonstrating the effectiveness of our Sensor-Language Alignment. The results shown are the averages from five random runs.
Impact of Alignment on HAR Performance Due to the limitations of Chronos’ pretraining sequence length, we were unable to use Chronos directly for HAR as a baseline for comparison. However, we tested the impact of bypassing Sensor-Language Alignment Stage by using Task-Aware Tuning Stage alone, directly perform HAR tasks using Chronos embeddings and the LLM. The results, shown in Figure 5, indicate that across all four datasets, SensorLLM outperforms the model that skips this step (Task-only). The performance of the Task-only model is either on par with or below the TS baselines, further highlighting the importance of our sensor-language alignment. These findings validate the effectiveness of aligning sensor-text modalities, as the integration of temporal and semantic features in the embeddings improves the LLM’s understanding of sensor data and significantly enhances HAR outcomes.
Table 2: The results for SensorLLM trained with/without text prompts. Task-only refers to conducting HAR directly bypassing sensor-language alignment. We report the F1-macro results in % with a mean and standard deviation of 5 repetitions.
Dataset Task-only (w/o prompts) Task-only (w/ prompts) SensorLLM (w/o prompts) SensorLLM (w/ prompts)
@@
Impact of Prompts To evaluate the contribution of additional textual information (statistical information for each channel’s sensor data) in Task-Aware Tuning Stage, we conducted experiments comparing SensorLLM’s performance with and without the inclusion of prompts. As shown in Table 2, the inclusion of prompts consistently improves F1-macro scores across all datasets, with a more pronounced effect when used in the full SensorLLM architecture. This indicates that the model effectively integrates both sensor and textual data, allowing it to better understand complex temporal patterns. The results highlight the advantages of leveraging multimodal inputs, which enrich the model’s representation of sensor data and lead to more accurate human activity recognition. Overall, the ability to process both sensor data and textual prompts not only boosts classification performance but also paves the way for LLMs to tackle more complex sensor-driven tasks in future applications.
Table 3: Cross-dataset experiments. We report the F1-macro results in % with a mean and standard deviation of 5 repetitions.
@@
Cross-Dataset Generalization To further assess the robustness of SensorLLM, we conducted cross-dataset experiments, where Sensor-Language Alignment Stage was trained on the USC-HAD dataset and Task-Aware Tuning Stage was trained on the UCI-HAR dataset (and vice versa). We chose these two datasets because they share the same sensor channels, although they differ in sample rates. As shown in Table 3, SensorLLM maintained strong performance, achieving results similar to those obtained when using the same dataset for both stages in Figure 4. These findings demonstrate that once the model effectively understands sensor data through modality alignment, retraining on new datasets for alignment becomes unnecessary. SensorLLM can directly perform downstream tasks across diverse datasets, laying the foundation for future TS-LLM models capable of generalizing across a wide range of time-series tasks.
7 Conclusion and future work
In this study, we introduced SensorLLM, a novel multimodal framework for aligning sensor data with automatically generated text. And unlike previous alignment methods which operate at the machine level, our approach aligns sensor data and text at the human perception level. SensorLLM thus can capture and interpret complex sensor data patterns—typically challenging for pre-trained LLMs—while outperforming or matching state-of-the-art models in HAR tasks. The results of cross-datasets experiments, without the need for dataset-specific alignment, also demonstrate SensorLLM’s strong generalizability and robustness. We have open-sourced our code and data generation methods, providing a valuable resource for the community and opening new avenues for aligning time-series and text data, particularly in domains with limited textual labels. Future work could explore more advanced analyses, such as chain-of-thought reasoning, sensor data generation, and zero-shot or few-shot learning, further enhancing LLMs’ potential in sensor-based tasks.

### v2

@@ 3.3 Task-Aware Tuning
As shown in Figure 2 (b), the Task-Aware Tuning stage refines the multimodal sensor-text embeddings for HAR. This stage integrates multi-channel sensor readings with activity labels, aligning temporal patterns with human activities. The input sensor data $\mathbf{X}$ is segmented into overlapping windows of size $L$ with a 50% overlap (Li et al., 2018), forming segments $\mathbf{X}_{S}\in\mathbb{R}^{S\times C\times L}$, where $S$ is the number of segments and $C$ is the number of channels. The pretrained alignment module from the first stage maps sensor data to activity labels, preserving inter-channel dependencies while learning activity-related patterns.
Input Embedding.
@@
For each sensor channel $c$, we retrieve its aligned sensor embeddings $\hat{o}_{s}^{c}$. These embeddings are then concatenated across all channels, along with their corresponding statistical features (mean and variance), to form the final input embedding:
$\hat{z}=\hat{o}_{s}^{1}\oplus\hat{o}_{s}^{2}\oplus\dots\oplus\hat{o}_{s}^{C}\oplus\hat{z}_{\text{stat}},$ (3)
where $\hat{z}_{\text{stat}}$ represents the statistical information, and $C$ is the number of channels. This ensures the model integrates both temporal and statistical characteristics for HAR.
Loss Function.
@@
The input token sequence is processed by the LLM, yielding a latent representation $\mathbf{H}\in\mathbb{R}^{K\times D}$, where $K$ is the number of tokens and $D$ is the embedding dimension. Due to causal masking, we extract the final hidden state, $\mathbf{h}=\mathbf{H}_{K}$, which encodes all preceding token information. This pooled vector is passed through a fully connected layer to produce a prediction vector of size $M$, where $M$ is the number of activity classes. The final class probabilities $\hat{y}_{i}$ are obtained via the softmax function, and the model is optimized using cross-entropy loss:
$\mathcal{L}_{cls}=-\sum_{i=0}^{M-1}y_{i}\log\hat{y}_{i},$ (4)
where $y_{i}$ is the ground truth label.
@@ 4 Experiments
In this section, we evaluate SensorLLM in enabling LLMs to interpret, reason about, and classify sensor data for HAR tasks. All experiments are conducted on NVIDIA A100-80G GPUs. To assess the LLM’s ability to learn and generalize from raw sensor inputs, we ensure that the same training and testing subjects are used in both the Sensor-Language Alignment and Task-Aware Tuning stages. This guarantees that test data in the second stage remains unseen during alignment, ensuring a fair evaluation of generalization. We select Chronos as the TS embedder because it has not been pre-trained on motion sensor data, making it an ideal candidate for evaluating our approach’s robustness in learning directly from unprocessed sensor signals.
4.1 Datasets
To evaluate the effectiveness and generalizability of SensorLLM, we conduct experiments on five publicly available HAR datasets: USC-HAD (Zhang and Sawchuk, 2012), UCI-HAR (Anguita et al., 2013), PAMAP2 (Reiss and Stricker, 2012), MHealth (Baños et al., 2014), and CAPTURE-24 (Chan et al., 2024). These datasets differ in sensor placement, sampling rates, channel configurations, and activity types, covering both controlled laboratory conditions and free-living environments. Additionally, they vary in scale, with USC-HAD, UCI-HAR, PAMAP2, and MHealth collected from a limited number of subjects, whereas CAPTURE-24 features a large-scale dataset with sensor recordings from 151 participants in real-world settings. Full dataset details, including subject count, sensor configurations, data splits, activity classes, preprocessing steps, and windowing strategies, are provided in Appendix A.6.
Metric USC-HAD UCI-HAR PAMAP2 MHealth CAPTURE-24
GPT-4o Ours GPT-4o Ours GPT-4o Ours GPT-4o Ours GPT-4o Ours
BLEU-1 41.43 57.68 37.97 56.78 46.35 60.20 49.97 61.38 46.58 57.10
ROUGE-1 54.92 68.32 51.24 67.63 58.08 69.92 61.11 71.20 58.21 68.11
ROUGE-L 49.00 64.17 44.88 63.05 50.30 66.25 51.99 67.83 48.88 60.90
METEOR 30.51 45.95 26.93 45.81 37.17 52.21 38.50 51.73 31.16 40.51
SBERT 77.22 86.09 76.05 85.01 82.71 87.31 83.15 86.66 83.11 84.83
SimCSE 86.96 93.09 90.23 92.51 89.64 93.82 92.10 93.38 90.10 92.20
GPT-4o 1.67 3.11 1.61 3.20 1.90 3.77 1.69 3.69 1.70 2.32
Human 2.10 4.16 1.94 4.04 2.38 4.70 1.74 4.56 2.30 3.10
@@ 4.2 Sensor Data Understanding
We train SensorLLM on all five datasets using the same following parameters in the Sensor-Language Alignment Stage: a learning rate of 2e-3, 8 epochs, batch size of 4, gradient accumulation steps of 8, and a maximum sequence length of 8192 for CAPTURE-24 and 4096 for the other datasets.
Evaluation Metrics.
@@
We assess SensorLLM’s ability to generate trend descriptions from sensor data, comparing it with the advanced GPT-4o 111gpt-4o-2024-08-06(OpenAI, 2024) to evaluate Sensor-Language Alignment. GPT-4o generates responses based on a predefined prompt (Appendix A.4), following our text template. We employ three evaluation methods:
@@
  • NLP Metrics. We measure surface-level similarity and n-gram overlap using BLEU-1 (Papineni et al., 2002), ROUGE-1, ROUGE-L (Lin, 2004), and METEOR (Banerjee and Lavie, 2005). We also use SBERT (Reimers and Gurevych, 2019) and SimCSE (Gao et al., 2021) to evaluate semantic similarity between model outputs and ground truth.
  • GPT-4o Evaluation. GPT-4o rates the generated trend descriptions on a scale of 1 to 5 (with 5 being the highest) by comparing each output to ground truth and providing explanatory feedback. As an advanced LLM, its evaluation ensures a semantic assessment of trend comprehension.
  • Human Evaluation. Five time-series experts (PhD students, postdocs, and academics) score accuracy and quality using the same criteria as GPT-4o, providing a human-centered perspective on the model’s outputs.
Appendix A.5 details all metrics and scoring criteria. We randomly sample 200 instances per dataset for both SensorLLM and GPT-4o, then average the results for comparison. Because reading and comparing lengthy sequences is difficult for human annotators, we conduct human evaluation on 20 shorter sequences per dataset (each containing at most 50 time steps).
Results.
@@
Table 1 compares SensorLLM and GPT-4o on the Sensor Data Trend Analysis tasks, showing that our model consistently outperforms GPT-4o across all metrics. BLEU-1, ROUGE-1, ROUGE-L, and METEOR primarily focus on surface-level lexical or n-gram overlaps and SBERT and SimCSE can capture factual correctness or deeper semantic similarities. Across all metrics, SensorLLM generates trend descriptions more closely aligned with the ground truth. GPT-4o evaluations further highlight SensorLLM’s superior ability to capture trend details and coherence, whereas GPT-4o struggles with complex numerical data and trend observations (Yehudai et al., 2024). Human evaluation also favors SensorLLM, particularly for shorter sequences. CAPTURE-24 results are weaker compared to other datasets, likely due to its longer sequences being trained with the same parameters. Overall, these findings validate the effectiveness of our Sensor-Language Alignment method in enhancing LLMs’ ability to interpret complex numerical sequence. Appendix A.9 provides qualitative examples of outputs from both models.
Method USC-HAD UCI-HAR PAMAP2 MHealth CAPTURE-24
PatchTST 45.2±$\pm 1.48$ 86.8±$\pm 0.84$ 82.0±$\pm 0.71$ 80.0±$\pm 1.58$ 35.6±$\pm 0.89$
Ns-Transformer 52.6±$\pm 2.30$ 88.0±$\pm 0.71$ 78.8±$\pm 0.84$ 77.2±$\pm 1.48$ 34.8±$\pm 1.10$
Informer 51.2±$\pm 1.30$ 86.6±$\pm 1.14$ 78.0±$\pm 1.58$ 74.0±$\pm 0.71$ 35.6±$\pm 0.55$
Transformer 49.6±$\pm 1.67$ 85.4±$\pm 0.89$ 77.0±$\pm 0.71$ 75.2±$\pm 1.30$ 32.8±$\pm 0.84$
iTransformer 48.4±$\pm 1.82$ 81.8±$\pm 0.84$ 76.6±$\pm 0.55$ 80.4±$\pm 1.14$ 19.8±$\pm 0.84$
TimesNet 52.2±$\pm 2.39$ 87.4±$\pm 1.14$ 76.2±$\pm 1.92$ 78.4±$\pm 1.52$ 34.8±$\pm 0.84$
GPT4TS 54.2±$\pm 2.05$ 88.2±$\pm 0.84$ 80.4±$\pm 0.89$ 76.4±$\pm 1.14$ 32.8±$\pm 1.10$
Chronos+MLP 44.2±$\pm 1.30$ 82.2±$\pm 0.84$ 79.8±$\pm 0.45$ 83.0±$\pm 0.71$ 38.0±$\pm 0.71$
DeepConvLSTM 48.8±$\pm 2.39$ 89.2±$\pm 0.84$ 78.4±$\pm 1.52$ 75.0±$\pm 1.87$ 40.4±$\pm 0.89$
DeepConvLSTMAtt 54.0±$\pm 2.12$ 89.6 ±$\pm 1.14$ 79.2±$\pm 1.30$ 77.4±$\pm 2.19$ 41.4±$\pm 0.55$
Attend 60.2±$\pm 2.17$ 93.2±$\pm 0.84$ 84.6±$\pm 1.14$ 83.4±$\pm 1.14$ 43.6±$\pm 0.55$
SensorLLM 61.2±$\pm 3.56$ 91.2±$\pm 1.48$ 86.2±$\pm 1.48$ 89.4±$\pm 3.85$ 48.6±$\pm 1.14$
Table 2: F1-macro results (%) for the Task-Aware Tuning Stage, presented as the mean and standard deviation over 5 random repetitions. The top results for each dataset are highlighted as follows: Bold for the best and underline for the second-best.
@@ 4.3 Human Activity Recognition
Setup.
In this section, we evaluate the performance of SensorLLM on HAR tasks. Each experiment runs for five trials, using 8 training epochs, a batch size of 4, gradient accumulation steps of 8, and a maximum sequence length of 4096. We report the F1 macro score A.8 to account for class imbalance across different activity categories.
Baselines.
We benchmark SensorLLM against 11 baselines across two categories: (i) TS models—Transformer (Vaswani et al., 2017), Informer (Zhou et al., 2021), NS-Transformer (Liu et al., 2022), PatchTST (Nie et al., 2023), TimesNet (Wu et al., 2023a), and iTransformer (Liu et al., 2024c); (ii) HAR models—DeepConvLSTM (Ordóñez and Roggen, 2016), DeepConvLSTMAttn (Murahari and Plötz, 2018), and Attend (Abedin et al., 2021). We also include Chronos+MLP and GPT4TS (Zhou et al., 2023a) for a more comprehensive comparison. Full baseline details are in Appendix A.7.
Results.
Table 2 reports the macro F1 scores (%) averaged over five random runs. SensorLLM achieves the highest performance on four datasets (USC-HAD, PAMAP2, MHealth, CAPTURE-24) and ranks second on UCI-HAR, demonstrating its effectiveness in handling diverse sensor data.
Notably, SensorLLM achieves a notable improvement on the CAPTURE-24 dataset, surpassing all baselines by a significant margin with a mean F1-macro score of 48.6%, which is 5.0% higher than Attend (43.6%). On USC-HAD, SensorLLM achieves the highest score of 61.2%, outperforming Attend, the second-best baseline, by 1.0%. Similarly, on PAMAP2, SensorLLM achieves a score of 86.2%, exceeding Attend (84.6%) by 1.6%. On MHealth, SensorLLM sets a new state-of-the-art with a score of 89.4%, surpassing Attend (83.4%) by 6.0%. These results highlight SensorLLM’s ability to consistently outperform existing methods across diverse datasets.
For UCI-HAR, SensorLLM achieves the second-best score (91.2%), slightly trailing Attend (93.2%). In contrast, Chronos+MLP shows only a slight improvement over iTransformer, the lowest-performing baseline (82.2% vs. 81.8%), indicating that Chronos embeddings alone have limited utility for HAR on this dataset. However, our framework significantly enhances their effectiveness, highlighting the robustness of our alignment approach.
@@ 5 Ablation Studies
Dataset Task-only SensorLLM
w/o prompts w/ prompts w/o prompts w/ prompts
@@
CAPTURE-24 44.8±$\pm 0.84$ 46.0±$\pm 0.71$ 47.2±$\pm 0.84$ 48.6±$\mathbf{\pm 1.14}$
Table 3: The results for SensorLLM trained with/without text prompts. Task-only refers to conducting HAR directly bypassing sensor-language alignment.
Impact of Alignment.
To evaluate the impact of alignment, we included the Chronos+MLP baseline in Section 4.3 to show that SensorLLM’s performance is not solely driven by Chronos embeddings. Additionally, we compared SensorLLM with a Task-only model, which skips the Sensor-Language Alignment Stage and directly applies Chronos embeddings and the LLM for HAR. As shown in Table 3, SensorLLM consistently outperforms the Task-only model across all five datasets, regardless of additional textual input. Notably, the Task-only model often performs on par with or worse than traditional TS baselines, underscoring the importance of our alignment method. These results confirm that Chronos embeddings alone are insufficient for optimal HAR performance and that alignment is essential for enabling LLMs to effectively understand sensor data.
Impact of Prompts.
To assess the role of additional textual information (e.g., statistical features for each sensor channel) in the Task-Aware Tuning Stage, we compared SensorLLM’s performance with and without prompts. As shown in Table 3, incorporating prompts consistently improves F1-macro scores across all datasets, with a more pronounced effect in the full SensorLLM architecture. This demonstrates that the model effectively integrates sensor and textual data, enhancing its ability to capture complex temporal patterns. The results highlight the benefits of multimodal inputs, which enrich sensor data representations and improve HAR accuracy. More broadly, the ability to process both sensor signals and textual prompts not only enhances classification performance but also extends LLMs’ potential for tackling complex sensor-driven tasks in future applications.
Figure 3: Effect of the number of alignment module layers.
Alignment Module Layers.
We investigate the effect of hidden layer depth in the alignment module MLP on UCI-HAR, PAMAP2, and MHealth. As shown in Figure 3, increasing the number of hidden layers from one ($1024\to 2048\to 4096$) to two ($1024\to 2048\to 3072\to 4096$) led to mixed results. The average F1-macro scores improved on UCI-HAR (91.2% → 92.0%) and MHealth (89.4% → 90.2%), but slightly dropped on PAMAP2 (86.2% → 85.8%). These results suggest that the optimal number of hidden layers varies by dataset.
Effect of Model Size on Performance.
To assess the impact of model size, we tested SensorLLM-3b, a resource-efficient variant using Chronos-base and Llama3.2-3b. Experiments were conducted on USC-HAD, UCI-HAR, and MHealth. As shown in Figure 4, SensorLLM-3b achieves slightly lower performance than SensorLLM-8b across all datasets, illustrating the trade-off between model size and accuracy. However, SensorLLM-3b remains competitive, outperforming Attend on USC-HAD and MHealth, while trailing it only on UCI-HAR. These results indicate that SensorLLM-3b offers a resource-efficient alternative, maintaining strong performance relative to other baselines.
Figure 4: Effect of Model Size.
Cross-Dataset Generalization.
To assess the robustness of SensorLLM, we conducted cross-dataset experiments, training the Sensor-Language Alignment Stage on USC-HAD and the Task-Aware Tuning Stage on UCI-HAR, and vice versa. These datasets share the same sensor channels but differ in sampling rates. As shown in Table 4, SensorLLM achieves performance comparable to training both stages on the same dataset. This indicates that once modality alignment is established, retraining on new datasets is unnecessary. SensorLLM effectively generalizes to downstream tasks across diverse datasets, demonstrating that the alignment stage enables the LLM to truly understand sensor data, rather than memorizing dataset-specific features. These results highlight SensorLLM’s potential for broad cross-dataset generalization, laying the groundwork for future TS-LLM models.
@@
Table 4: Cross-dataset experiments.
6 Conclusions
We introduced SensorLLM, a multimodal framework that aligns sensor data with automatically generated text at a human-perception level, moving beyond machine-level alignment. SensorLLM effectively captures complex sensor patterns, achieving superior performance in HAR tasks. Experiments across diverse datasets demonstrate its robustness in handling variable-length sequences, multi-channel inputs, and textual metadata. Cross-dataset results further highlight its strong generalizability without requiring dataset-specific alignment. This work establishes a foundation for Sensor-Text MLLMs, with potential applications for sensor data analysis. We release our code and data generation pipeline to facilitate future research on integrating time-series and text, particularly in low-resource domains.
7 Limitations
While SensorLLM demonstrates strong performance in aligning sensor data with LLMs, certain limitations remain, offering directions for future exploration.
Classifier-Based Design.
To ensure fair comparisons with existing HAR models (classification models), we adopt a classifier for downstream tasks rather than fully exploiting the LLM’s generative abilities. Although our results verify that the Sensor-Language Alignment Stage can generalize across datasets, relying on a fixed-class classifier may constrain the model’s adaptability to new activity categories. Future work could explore using the LLM in a generative or prompt-based capacity, enabling broader application scenarios such as activity discovery or open-set recognition.
Scope of Sensor-Text Alignment.
Our alignment focuses on mapping sensor data to trend-descriptive text, demonstrating clear benefits for LLM-based HAR. However, human-intuitive descriptions of sensor data extend beyond trend changes—incorporating frequency-domain features, periodicity, and higher-order patterns may further enhance an LLM’s ability to interpret time-series data. Future research could investigate whether aligning text with alternative sensor characteristics improves time-series reasoning. This could expand the potential of multimodal NLP applications in sensor-driven tasks beyond activity recognition.
8 Acknowledgements
This research includes computations using the computational cluster Wolfpack supported by School of Computer Science and Engineering at UNSW Sydney.